Find the distance between points (-1,-1)and (2,5)
I am really not sure what I did wrong on this one. It was a multiple choice question. A is 6 square root of 5. B is 3 square root of 6. C is 5 square root of 6. And D is 6 square root of 3. I came up with C. 5 square root of 6. Please tell me what I did wrong....

Found 2 solutions by stanbon, checkley71:
Answer by stanbon(48568)

Find the distance between points (-1,-1)and (2,5)
-------------
distance = sqrt[(5--1)^2 + (2--1)^2]
distance = sqrt[36 + 9]
distance = sqrt[45]
distance = 3sqrt5
===================
Cheers,
Stan H.

Answer by checkley71(8403)

I am really not sure what I did wrong on this one. It was a multiple choice question. A is 6 square root of 5. B is 3 square root of 6. C is 5 square root of 6. And D is 6 square root of 3. I came up with C. 5 square root of 6. Please tell me what I did wrong....
DON'T KNOW WHAT YOU DID TO SOLVE THIS.
YOU COULD HAVE PLOTTED THESE 2 POINTS ON GRAPH PAPER, CONSTRUCTED THE RIGHT TRIANGLE & THEN SOLVED FOR THE HYPOTENUSE. USE INCHES FOR THE 2 POINTS SO YOU'LL BE MORE ACCURATE ON THE FINAL MEASUREMENT.
MEASURING WOULD NOT HAVE HELPED AS YOU ONLY HAVE SQRT IN THE ANSWERS.
HOWEVER CONVERTING THESE ANSWERS TO NUMBERS I GET:
6SQRT5=6*2.23=13.42 WHICH IS NOT THE ANSWER.
3SQRT6=3*2.45=7.35 WHICH IS NOT THE ANSWER.
5SQRT6=5*2.45=14.7 WHICH IS NOT THE ANSWER.
6SQRT3=6*1.732=10.39 WHICH IS NOT THE ANSWER.
CHECK THESE VALUES OUT BY MEASURING THE HYPOTENUSE ON THE GRAPH PAPER.
-----------------------------------------------------
FIRST WE FIND THE X DISTANCE: 2+1=3
NOW WE FIND THE Y DISTANCE: 5+1=6
NOW WE HAVE TWO SIDES OF A RIGHT TRIANGLE WITH THE DISTANCE BETWEEN THEM IS THE HYPOTENUSE OF THIS RIGHT TRIANGLE:
USING THE PYTHAGOREON THEORUM A^2+B^2=C^2 WE FIND THE HYPOTENUSE (C).
3^2+6^2=C^2
9+36=C^2
45=C^2
C=SQRT45
C=6.7 IS THE LENGTH OF THE HYPOTENUSE.